# Exploring cosmological imprints of phantom crossing with dynamical dark energy in Horndeski gravity

### v1

@@ I Introduction
Advances in observational techniques and statistical inference have ushered cosmology into a precision era, enabling constraints on cosmological parameters at the sub-percent level. While this progress has strengthened confidence in the standard Lambda-Cold Dark Matter ($\Lambda$CDM) cosmology, the past decade has also revealed inconsistencies among certain cosmological parameters inferred from different datasets. The most prominent of these are the Hubble tension and the growth tension. Hubble tension refers to the significant discrepancy between the Hubble constant, $H_{0}$, measured from local universe observations, such as distance ladder calibrations, and the value inferred from Cosmic Microwave Background (CMB) observations under the assumption of the $\Lambda$CDM model. For instance, the SH0ES collaboration Riess et al. (2022) reports a value of $H_{0}=73.04\pm 1.04$ km s-1 Mpc-1, while the Planck best-fit $\Lambda$CDM value is $H_{0}=67.4\pm 0.5$ km s-1 Mpc-1, resulting in a tension of $5\sigma$ significance Aghanim et al. (2020a). The growth tension pertains to discrepancies in the late-time cosmic structure formation parameters $\sigma_{8}$ (the amplitude of matter density fluctuations averaged over a scale of $8h^{-1}$ Mpc) and $S_{8}\equiv\sigma_{8}\sqrt{\Omega_{m}/0.3}$, where $\Omega_{m}$ is the present value of the matter density parameter. Weak lensing surveys like the Kilo-Degree Survey (KiDS) Asgari et al. (2021) report $S_{8}=0.759\pm 0.021$, which is lower than the CMB-inferred value of $S_{8}=0.834\pm 0.016$, leading to a $2-3\sigma$ tension. While this level of tension might seem modest, it can introduce significant uncertainty into cosmological simulations of large-scale structures. Moreover, the growth tension can also be linked to deviations from General Relativity (GR) when the growth rate of matter perturbations is parameterized using the ’growth index’ $\gamma$. The growth rate is approximated as $f(a)=\Omega_{m}(a)^{\gamma}$, where $\gamma=0.55$ is the GR prediction for a flat $\Lambda$CDM model. However, as highlighted in Nguyen et al. (2023), a model-independent analysis incorporating $f\sigma_{8}$ measurements from peculiar velocities and Redshift Space Distortions (RSD), combined with Planck 2018 data, finds $\gamma=0.639^{+0.025}_{-0.024}$, a $3.7\sigma$ deviation from the GR expectation. This elevated $\gamma$ suggests a suppressed growth of structures during the dark energy-dominated era, aligning with the lower $S_{8}$ value inferred from weak lensing measurement and hinting at a possible deviation from GR. Assuming that these tensions are not due to some unaccounted observational systematics in the analysis, several approaches have been proposed to resolve them by modifying the standard $\Lambda$CDM model. In the case of the Hubble tension, solutions fall into two broad categories: modifying the physics before recombination to reduce the sound horizon or introducing late-time modifications to the expansion history. Both approaches aim to preserve the angular size of the sound horizon at the time of recombination, $\theta_{*}$, a precisely measured cosmological parameter from CMB observations (See, for instance, Refs. Knox and Millea (2020); Schöneberg et al. (2022); Di Valentino et al. (2021); Abdalla et al. (2022); Hu and Wang (2023) for a comprehensive review). However, most of the solutions that attempt to resolve the Hubble tension exacerbate the growth tension. In Refs. Jedamzik et al. (2021); Vagnozzi (2023), it has been demonstrated that early-universe solutions alone are insufficient to alleviate both tensions, often requiring a combination of early- and late-time modifications (see Simon et al. (2024), for example). Meanwhile, simple late-time modifications, such as a phantom dark energy equation of state, struggle to reconcile with Baryon Acoustic Oscillation (BAO) measurements and Supernova data, which tightly constrain the late-time expansion history Camarena and Marra (2021); Yang et al. (2021). It has also been pointed out in Heisenberg et al. (2023); Tutusaus et al. (2023) that for late-time modifications to resolve cosmological tensions without disturbing the CMB measurements, it requires a phantom crossing, where the dark energy equation must transit between phantom and non-phantom regimes at some redshift. As a result, the simultaneous resolution of both the $H_{0}$ and $S_{8}$ tensions remains a complex and challenging problem.
@@
In recent years, with the influx of low-redshift data, parametric and non-parametric reconstructions of the expansion history and dark energy properties, such as energy density and the equation of state, have been pursued to better understand cosmological tensions and the nature of dark energy. Many of these studies have pointed out that the late time expansion history may not be as trivial as explained by the cosmological constant but rather could be more dynamical in nature with interesting features like phantom crossing or the negative dark energy density at high redshifts Wang et al. (2018); Dutta et al. (2020); Escamilla and Vazquez (2023); Akarsu et al. (2023a); Malekjani et al. (2024); Gómez-Valent et al. (2024); Sabogal et al. (2024). The concept of negative dark energy at high redshifts was initially explored as a potential explanation for the anomalous Ly-$\alpha$ measurements of the expansion history by eBOSS around $z\sim 2.33$, which could not be explained by the $\Lambda$CDM model du Mas des Bourboux et al. (2020); Aubourg et al. (2015); Sahni et al. (2014); Manoharan (2024). Recent studies have examined scenarios wherein dark energy density or equation of state is parametrized to accommodate these features, and they are found to give a good fit to the observational data, reconciling cosmological tensions Akarsu et al. (2020); Sen et al. (2022); Akarsu et al. (2023b); Adil et al. (2024). Interestingly, the first data release of BAO measurements by Dark Energy Spectroscopic Instrument (DESI) has also hinted towards a dynamical nature of dark energy when combined with CMB and Supernovae data Adame et al. (2024a, b); Calderon et al. (2024); Lodha et al. (2024). This analysis is conducted for a CPL parameterization of the dark energy equation of state Chevallier and Polarski (2001), i.e., specifically within the $w_{0}w_{a}$ scenario. Subsequent reconstructions of the equation of state, using DESI, CMB, and SNIa data, have confirmed that this dynamical behaviour is not merely a consequence of the chosen CPL parameterization but is instead a feature supported by the data itself Ye et al. (2024); Giarè et al. (2024); Mukherjee and Sen (2024); Chudaykin and Kunz (2024). Notably, these studies have also uncovered hints of a phantom crossing at low redshifts, where the equation of state transitions to $w<-1$ around $z<1$. This makes it an intriguing direction to explore and revive our understanding of dark energy. If future observations continue to support evidence of its dynamical nature, along with the possibility of phantom behavior at certain redshifts, such phenomena would not be explained within the GR framework without invoking instabilities. Henceforth, it calls for the need to go beyond GR. Horndeski theory, a generalized scalar-tensor theory of gravity, is an excellent avenue to explore the simplest extension of GR, offering a rich phenomenology Horndeski (1974); Kobayashi (2019); Bellini and Sawicki (2014). With a Lagrangian constructed from a scalar field and metric tensor, the theory gives second-order equations of motion. In fact, Horndeski theory can accommodate various subclasses of Modified Gravity (MG) models, including Brans-Dicke, Galileon, Quintessence, K-essence, and even interacting dark sectors Bansal et al. (2024) (such as chameleons), etc., along with GR as well. What makes Horndeski theory interesting is its phenomenological framework for exploring dynamical dark energy scenarios, particularly in light of emerging cosmological tensions. This is because it can accommodate novel features like phantom divide without ghost or gradient stabilities, as shown in various studies, which have become even more intriguing after the DESI data release Matsumoto (2018); Ye et al. (2024); Chudaykin and Kunz (2024).
@@ III The model
To investigate the evolution of both the background and perturbations, we implement our model in the hi_class code. The governing equations for the background dynamics—including the Friedmann equation, the scalar field evolution, and its corresponding energy density and pressure—are derived from the general expressions provided in Appendix A, where we substitute Eq. (4) to specialize them for our scenario. In particular, we adopt a linear potential for the scalar field, $V(\phi)=V_{0}\phi$. The hi_class code evolves the background equations by using initial conditions on Hubble parameter, scalar field and its derivative while $V_{0}$ is used as a tuning parameter. This parameter is calibrated via the shooting method to ensure the correct present-day value of the dark energy density. Consequently, the model contains three free parameters— $c_{1}$,$c_{2}$ and $c_{3}$. Varying these parameters reveals rich and interesting features in the dynamics of the scalar field and the expansion history of the universe. In addition to the background evolution, hi_class computes the evolution of key cosmological observables, such as the growth rate, the matter power spectrum, and the CMB power spectrum. This allows for a comprehensive exploration of how different parameter choices impact both large-scale structure formation and the CMB anisotropies. Notably, the setup is designed such that when the model parameters are set to zero ($c_{1}=c_{2}=c_{3}=0$), the system reverts to canonical Quintessence, which is similar to $\Lambda$CDM universe, ensuring consistency with the standard cosmological model as a baseline. Finally, to ensure a consistent and physically viable model, the evolution is kept free from ghost and gradient instabilities by using appropriate configurations within the hi_class code. In hi_class, a mix of natural units ($\hbar=c=1$) and cosmological conventions (expressing all quantities in terms of $\mathrm{Mpc}$) is used, with the scalar field normalized to be dimensionless. In this framework, $c_{1}$ and $c_{3}$ are dimensionless parameters, while $c_{2}$ is expressed in units of $\mathrm{Mpc}^{2}$.
@@ IV.1 Background
The background dynamics is governed by the homogeneous evolution of the energy density and pressure through the Friedmann equation. In our model, the scalar field serves as the dark energy component. Its effect on the background dynamics is most prominently visible in the evolution of the Hubble parameter and the dark energy equation of state. To maintain consistency with the physics of the early universe, we set initial conditions for the scalar field shortly after recombination, ensuring alignment with the minimal quintessence model, where dark energy density remains negligible. Consequently, the Hubble parameter initially follows the standard $\Lambda$CDM evolution and diverges at later times as the strength of various coupling parameters comes into play. Figure 1 illustrates the impact of these parameters on the Hubble expansion rate, with all other cosmological parameters fixed to the Planck 2018 best-fit values Aghanim et al. (2020a).
@@
For non-zero and positive values of $c_{3}$, energy density of the scalar field reaches negative values at high redshifts due to the dominance of the third term in Eq. (5). One straightforward consequence is that this reduces the expansion rate in comparison to $\Lambda$CDM during the period where energy density remains negative. Later on, as the universe expands and $H(z)$ decreases, other terms become dominant, leading to a positive energy density for the scalar field. This transition to a positive dark energy density is essential to explain the observed accelerated expansion at low redshifts. This transition from negative to positive energy density involves crossing zero, resulting in a singularity in the equation of state, which appears as a pronounced spike in numerical calculations, as shown in Figure 2. The precise location of these singularities is also influenced by the strength of the self-interaction terms, $c_{1}$ and $c_{2}$. Throughout this process, the scalar field consistently maintains negative pressure, allowing the equation of state to take on both positive and negative values across different cosmological epochs as the energy density evolves. The negative energy density at high redshifts is not problematic, as its effect remains subdominant during these epochs, and the total energy density is always positive. In the absence of non-minimal coupling (i.e., for $c_{3}=0$), no singularities or negative energy densities occur, as depicted by the purple and brown curves. Finally, with a combination of self-interactions and non-minimal coupling, we see interesting features in the dynamics where the former mainly controls the very late evolution while later controlling relatively high redshifts.
@@ Appendix A Horndeski Theory
In our analysis, we ensure that the stability conditions given in Eq. (A19) are consistently satisfied. This imposes stringent constraints on the parameter space of $c_{1}$, $c_{2}$ and $c_{3}$. Notably, the parameter $c_{2}$ plays a pivotal role in maintaining $Q_{s}>0$, thus preventing ghost instabilities while allowing the equation of state $w_{\phi}$ to enter the phantom regime. Another distinctive feature of Horndeski theories is the time evolution of the speed of tensor perturbations, given byDe Felice and Tsujikawa (2012),

### v2

@@ I Introduction
Advances in observational techniques and statistical inference have ushered cosmology into a precision era, enabling constraints on cosmological parameters at the sub-percent level. While this progress has strengthened confidence in the standard Lambda-Cold Dark Matter ($\Lambda$CDM) cosmology, the past decade has also revealed inconsistencies among certain cosmological parameters inferred from different datasets. The most prominent of these are the Hubble tension and the growth tension. Hubble tension refers to the significant discrepancy between the Hubble constant, $H_{0}$, measured from local universe observations, such as distance ladder calibrations, and the value inferred from Cosmic Microwave Background (CMB) observations under the assumption of the $\Lambda$CDM model. For instance, the SH0ES collaboration Riess et al. (2022) reports a value of $H_{0}=73.04\pm 1.04$ km s-1 Mpc-1, while the Planck best-fit $\Lambda$CDM value is $H_{0}=67.4\pm 0.5$ km s-1 Mpc-1, resulting in a tension of $5\sigma$ significance Aghanim et al. (2020a). The growth tension pertains to discrepancies in the late-time cosmic structure formation parameters $\sigma_{8}$ (the amplitude of matter density fluctuations averaged over a scale of $8h^{-1}$ Mpc) and $S_{8}\equiv\sigma_{8}\sqrt{\Omega_{m}/0.3}$, where $\Omega_{m}$ is the present value of the matter density parameter. Weak lensing surveys like the Kilo-Degree Survey (KiDS) Asgari et al. (2021) report $S_{8}=0.759\pm 0.021$, which is lower than the CMB-inferred value of $S_{8}=0.834\pm 0.016$, leading to a $2-3\sigma$ tension. While this level of tension might seem modest, it can introduce significant uncertainty into cosmological simulations of large-scale structures. Moreover, the growth tension can also be linked to deviations from General Relativity (GR) when the growth rate of matter perturbations is parameterized using the ’growth index’ $\gamma$. The growth rate is approximated as $f(a)=\Omega_{m}(a)^{\gamma}$, where $\gamma=0.55$ is the GR prediction for a flat $\Lambda$CDM model. However, as highlighted in Nguyen et al. (2023), a model-independent analysis incorporating $f\sigma_{8}$ measurements from peculiar velocities and Redshift Space Distortions (RSD), combined with Planck 2018 data, finds $\gamma=0.639^{+0.025}_{-0.024}$, a $3.7\sigma$ deviation from the GR expectation. This elevated $\gamma$ suggests a suppressed growth of structures during the dark energy-dominated era, aligning with the lower $S_{8}$ value inferred from weak lensing measurement and hinting at a possible deviation from GR. Assuming that these tensions are not due to some unaccounted observational systematics in the analysis, several approaches have been proposed to resolve them by modifying the standard $\Lambda$CDM model. In the case of the Hubble tension, solutions fall into two broad categories: modifying the physics before recombination to reduce the sound horizon or introducing late-time modifications to the expansion history. Both approaches aim to preserve the angular size of the sound horizon at the time of recombination, $\theta_{*}$, a precisely measured cosmological parameter from CMB observations (See, for instance, Refs. Knox and Millea (2020); Schöneberg et al. (2022); Di Valentino et al. (2021); Abdalla et al. (2022); Hu and Wang (2023) for a comprehensive review). However, most of the solutions that attempt to resolve the Hubble tension exacerbate the growth tension. In Refs. Jedamzik et al. (2021); Vagnozzi (2023), it has been demonstrated that early-universe solutions alone are insufficient to alleviate both tensions, often requiring a combination of early- and late-time modifications (see Simon et al. (2024), for example). Meanwhile, simple late-time modifications, such as a phantom dark energy equation of state, struggle to reconcile with Baryon Acoustic Oscillation (BAO) measurements and Supernova data, which tightly constrain the late-time expansion history Camarena and Marra (2021); Yang et al. (2021). It has also been pointed out in Heisenberg et al. (2023); Tutusaus et al. (2023); Lee et al. (2022) that for late-time modifications to resolve cosmological tensions without disturbing the CMB measurements, it requires a phantom crossing, where the dark energy equation must transit between phantom and non-phantom regimes at some redshift. As a result, the simultaneous resolution of both the $H_{0}$ and $S_{8}$ tensions remains a complex and challenging problem.
@@
In recent years, with the influx of low-redshift data, parametric and non-parametric reconstructions of the expansion history and dark energy properties, such as energy density and the equation of state, have been pursued to better understand cosmological tensions and the nature of dark energy. Many of these studies have pointed out that the late time expansion history may not be as trivial as explained by the cosmological constant but rather could be more dynamical in nature with interesting features like phantom crossing or the negative dark energy density at high redshifts Wang et al. (2018); Dutta et al. (2020); Escamilla and Vazquez (2023); Akarsu et al. (2023a); Malekjani et al. (2024); Gómez-Valent et al. (2024); Sabogal et al. (2024). The concept of negative dark energy at high redshifts was initially explored as a potential explanation for the anomalous Ly-$\alpha$ measurements of the expansion history by eBOSS around $z\sim 2.33$, which could not be explained by the $\Lambda$CDM model du Mas des Bourboux et al. (2020); Aubourg et al. (2015); Sahni et al. (2014); Manoharan (2024). Recent studies have examined scenarios wherein dark energy density or equation of state is parametrized to accommodate these features, and they are found to give a good fit to the observational data, reconciling cosmological tensions Akarsu et al. (2020); Teng et al. (2021); Sen et al. (2022); Akarsu et al. (2023b); Adil et al. (2024); Dwivedi and Högås (2024). Interestingly, the first data release of BAO measurements by Dark Energy Spectroscopic Instrument (DESI) has also hinted towards a dynamical nature of dark energy when combined with CMB and Supernovae data Adame et al. (2024a, b); Calderon et al. (2024); Lodha et al. (2024). This analysis is conducted for a CPL parameterization of the dark energy equation of state Chevallier and Polarski (2001), i.e., specifically within the $w_{0}w_{a}$ scenario. Subsequent reconstructions of the equation of state, using DESI, CMB, and SNIa data, have confirmed that this dynamical behaviour is not merely a consequence of the chosen CPL parameterization but is instead a feature supported by the data itself Ye et al. (2024); Giarè et al. (2024); Mukherjee and Sen (2024); Chudaykin and Kunz (2024). Notably, these studies have also uncovered hints of a phantom crossing at low redshifts, where the equation of state transitions to $w<-1$ around $z<1$. This makes it an intriguing direction to explore and revive our understanding of dark energy. If future observations continue to support evidence of its dynamical nature, along with the possibility of phantom behavior at certain redshifts, such phenomena would not be explained within the GR framework without invoking instabilities Vikman (2005); Deffayet et al. (2010); Wolf et al. (2024). Henceforth, it calls for the need to go beyond GR. Horndeski theory, a generalized scalar-tensor theory of gravity, is an excellent avenue to explore the simplest extension of GR, offering a rich phenomenology Horndeski (1974); Kobayashi (2019); Bellini and Sawicki (2014). With a Lagrangian constructed from a scalar field and metric tensor, the theory gives second-order equations of motion. In fact, Horndeski theory can accommodate various subclasses of Modified Gravity (MG) models, including Brans-Dicke, Galileon, Quintessence, K-essence, and even interacting dark sectors Bansal et al. (2024) (such as chameleons), etc., along with GR as well. What makes Horndeski theory interesting is its phenomenological framework for exploring dynamical dark energy scenarios, particularly in light of emerging cosmological tensions. This is because it can accommodate novel features like phantom divide without ghost or gradient stabilities, as shown in various studies, which have become even more intriguing after the DESI data release Matsumoto (2018); Ye et al. (2024); Chudaykin and Kunz (2024).
@@ III The model
To investigate the evolution of both the background and perturbations, we implement our model in the hi_class code. The governing equations for the background dynamics—including the Friedmann equation, the scalar field evolution, and its corresponding energy density and pressure—are derived from the general expressions provided in Appendix A, where we substitute Eq. (4) to specialize them for our scenario. In particular, we adopt a linear potential for the scalar field, $V(\phi)=V_{0}\phi$. The hi_class code evolves the background equations by using initial conditions on Hubble parameter, scalar field and its derivative while $V_{0}$ is used as a tuning parameter. This parameter is calibrated via the shooting method to ensure the correct present-day value of the dark energy density. The initial conditions for the scalar field and its derivative are set shortly after recombination at $z\sim 1000$ with values $\phi_{i}=10.0$ and $\phi^{{}^{\prime}}_{i}=10^{-10}$ in Planck units where $\phi^{{}^{\prime}}=d\phi/d\tau$, where $\tau$ is the conformal time. Consequently, the model contains three free parameters— $c_{1}$,$c_{2}$ and $c_{3}$. The choice of initial conditions on $\phi$ and $\phi^{{}^{\prime}}$ is motivated such that when the model parameters are set to zero ($c_{1}=c_{2}=c_{3}=0$), the system reverts to canonical Quintessence, which is similar to $\Lambda$CDM universe, ensuring consistency with the standard cosmological model as a baseline. Varying the free parameters reveals rich and interesting features in the dynamics of the scalar field and the expansion history of the universe. In addition to the background evolution, hi_class computes the evolution of key cosmological observables, such as the growth rate, the matter power spectrum, and the CMB power spectrum. This allows for a comprehensive exploration of how different parameter choices impact both large-scale structure formation and the CMB anisotropies. Finally, to ensure a consistent and physically viable model, the evolution is kept free from ghost and gradient instabilities by using appropriate configurations within the hi_class code. In hi_class, a mix of natural units ($\hbar=c=1$) and cosmological conventions (expressing all quantities in terms of $\mathrm{Mpc}$) is used, with the scalar field normalized to be dimensionless. In this framework, $c_{1}$ and $c_{3}$ are dimensionless parameters, while $c_{2}$ is expressed in units of $\mathrm{Mpc}^{2}$.
@@ IV.1 Background
The background dynamics is governed by the homogeneous evolution of the energy density and pressure through the Friedmann equation. In our model, the scalar field serves as the dark energy component. Its effect on the background dynamics is most prominently visible in the evolution of the Hubble parameter and the dark energy equation of state. To maintain consistency with the physics of the early universe, we set initial conditions for the scalar field shortly after recombination, ensuring alignment with the minimal quintessence model, where dark energy density remains negligible. Consequently, the Hubble parameter initially follows the standard $\Lambda$CDM evolution and diverges at later times as the strength of various coupling parameters comes into play. Figure 1 illustrates the impact of these parameters on the Hubble expansion rate and thereby on the $H_{0}$, with all other cosmological parameters fixed to the Planck 2018 best-fit values Aghanim et al. (2020a).
@@
For non-zero and positive values of $c_{3}$, energy density of the scalar field reaches negative values at high redshifts due to the dominance of the third term in Eq. (4). One straightforward consequence is that this reduces the expansion rate in comparison to $\Lambda$CDM during the period where energy density remains negative. Later on, as the universe expands and $H(z)$ decreases, other terms become dominant, leading to a positive energy density for the scalar field. Despite the negative values of $\rho_{\phi}$ at high redshifts, the total energy density of the universe remains positive, as the dark energy contribution is subdominant during those epochs. Consequently, the entire cosmological evolution remains free of instabilities. However, the transition to a positive dark energy density is essential to explain the observed accelerated expansion at low redshifts. This transition from negative to positive energy density involves crossing zero, resulting in a singularity in the equation of state, which appears as a pronounced spike due to finite step size in numerical calculations, as shown in Figure 2. This is physically acceptable, as both energy density and pressure contribute similarly to gravitational effects, and the absence of one does not inherently lead to any issues. Since $w$ denotes a derived quantity used for convenience to describe the behavior of dark energy, a singularity does not cause any unphysical behavior in the underlying dynamics. The precise location of these singularities is also influenced by the strength of the self-interaction terms, $c_{1}$ and $c_{2}$. Throughout this process, the scalar field consistently maintains negative pressure, allowing the equation of state to take on both positive and negative values across different cosmological epochs as the energy density evolves. In the absence of non-minimal coupling (i.e., for $c_{3}=0$), no singularities or negative energy densities occur, as depicted by the purple and brown curves. In summary, the dynamics reveal intriguing features, with self-interactions ($c_{1}$ and $c_{2}$) primarily governing the late-time evolution of $H(z)$, while the non-minimal coupling ($c_{3}$) dictates the behavior of scalar field at high redshifts. It is the interplay between $c_{1}$, $c_{2}$, and $c_{3}$ that significantly influences deviations from $\Lambda$CDM, resulting in higher or lower $H_{0}$ values. However, the optimal values for these parameters, along with the cosmological parameters, which give a good fit to the observational data, are determined through detailed data analysis, as discussed in Sec. V.
@@ Appendix A Horndeski Theory
In our analysis, we ensure that the stability conditions given in Eq. (A19) are consistently satisfied. This imposes stringent constraints on the parameter space of $c_{1}$, $c_{2}$ and $c_{3}$. Notably, the parameter $c_{2}$ plays a pivotal role in maintaining $Q_{s}>0$, thus preventing ghost instabilities while allowing the equation of state $w_{\phi}$ to enter the phantom regime. Another distinctive feature of Horndeski theories is the time evolution of the speed of tensor perturbations, given by De Felice and Tsujikawa (2012); Tiwari et al. (2023),